For each structure shown below, determine the numb
Answer #1
Concepts and reason

The concept used in solving the given problem is the Huckel’s rule and estimation of lone pairs and the π{\rm{\pi }} bonds. The presence of π{\rm{\pi }} -bonds and lone pair of the electrons helps in determining the number of electrons and the Huckel’s rule help in the determination of the aromaticity of the compound.

Fundamentals

A π{\rm{\pi }} -bond consists of 22 π{\rm{\pi }} -electrons and two electrons are present in a lone pair of electrons.

• Aromatic compounds:

1.It should follow Huckel’s rule, that is, the compound should contain (4n+2)π\left( {4n + 2} \right)\pi electrons in cyclic conjugation where is a whole number.

2.It should be planar.

3.It should be cyclic.

4.It should have continuous delocalization of electrons.

• Antiaromatic compounds do not follow (4n+2)π\left( {4n + 2} \right)\pi electrons instead they have (4n)π\left( {4n} \right)\pi electrons in cyclic conjugation.

• Non-aromatic compounds are the one which are not planar and neither belongs to aromatic nor antiaromatic compounds.

The number of π{\rm{\pi }} (pi) electrons present in the structure given below is 44 .

Furthermore, the given compound is antiaromatic in nature.

The number of π{\rm{\pi }} -electrons present in the structure given below is 66 .

Furthermore, the given compound is aromatic in nature.

The number of π{\rm{\pi }} -electrons present in the structure given below is 66 .

Furthermore, the given compound is aromatic in nature.

Ans:

The compound mentioned below has 44 π{\rm{\pi }} -electrons present and is nonaromatic in nature.

The compound mentioned below has 66 π{\rm{\pi }} -electrons present and is aromatic in nature.

The compound mentioned below has 66 π{\rm{\pi }} -electrons present and is aromatic in nature.